OpenStudy (anonymous):

NEED Help ? anybody? A manufacturing unit currently operates at 80 percent of its capacity. The profit function for the unit at the optimum output, x, is given by p(x) = -0.1x^2 + 80x − 60. If the function f(x) models the current capacity of the unit, the composite function giving the unit's current profit function is a.f(p(x))=-0.064x^2+6.4x-60 b.p(f(x))=-0.64x^2+0.64x-60 c.p(f(x))=-0.064x^2+64x-60 d.p(f(x))=-0.64x^2+6.4x-60 e.f(p(x))=-0.064x^2+64x-60 If the optimum output is 500 units, the current profit is $a.15,400 b.15,940 c.16,060 d.16,600

OpenStudy (jdoe0001):

we seem to have a missing f(x) for one

OpenStudy (jdoe0001):

f( px ) is the composite of both functions, however we only have f(x)

OpenStudy (jdoe0001):

( px ) is the composite of both functions, however we only have p(x) only rather

OpenStudy (anonymous):

i send a screenshoot so you can see

OpenStudy (anonymous):

the options for the first are a. f(p(x))=-0.064x^2+6.4x-60 b. p(f(x))=-0.64x^2+0.64x-60 c. p(f(x))=-0.064x^2+64x-60 d. p(f(x))=-0.64x^2+6.4x-60 e. f(p(x))=-0.064x^2+64x-60